OpenStudy (anonymous):

I think I have the answer, can I get a second opinion? "An equation is shown below: 4(2x - 5) + 15 = 11 Write the steps you will use to solve the equation and explain each step." I got x = 2, but I did mental math and I'm not really good at writing it down. Can someone help me word it?

OpenStudy (lynfran):

multiply out the bracket step 1

OpenStudy (welshfella):

4(2x - 5) + 15 = 11 distribute 4 over the parentheses 8x - 20 + 15 = 11 simplify 8x - 5 = 11 add 5 to both sides 8x = 11 + 5 = 16 divide both sides by 8 x = 16/8 = 2

OpenStudy (lynfran):

yep @welshfella

OpenStudy (lynfran):

and ur answer was correct...@flowury

OpenStudy (anonymous):

Aaa thank you guys so much. :)